1. Solve the equation below for d, in terms of a,b,and c.
a/b = c/d
2. Solve the equation below for c, in terms of a,b,and d.
a/b = c/d

3. Solve the equation below for m, in terms of y, x,and b.
y=mx+b

4. Solve the equation below for a, in terms of b and c.
a+4b-3c=10

5. Solve the equation below for y, in terms of x.
2x+5y=3

6. Solve the equation below for I in terms of v, w, and h.
v= i w h

7. Solve the equation below for x, in terms of y, x,and b.
y=mx+b

8. Solve the equation below for c, in terms of a and b.
a+4b-3c=10




Answer by Osran_Shri(18)

1. Solve the equation below for d, in terms of a,b,and c.
a/b = c/d
a/b = c/d multiply by d on both sides to get
d(a/b) = (c/d)d
da / b = c
d = (c * b) / a
2. Solve the equation below for c, in terms of a,b,and d.
a/b = c/d
c = (a * d) / b
3. Solve the equation below for m, in terms of y, x,and b.
y=mx+b
mx = y-b
m = (y-b)/x
4. Solve the equation below for a, in terms of b and c.
a+4b-3c=10
a = 10-4b+3c
5. Solve the equation below for y, in terms of x.
2x+5y=3
5y = 3-2x y = (3-2x)/5
6. Solve the equation below for I in terms of v, w, and h.
v= i w h
i = v / (w*h)
7. Solve the equation below for x, in terms of y, x,and b.
y=mx+b
mx = y-b
x = (y-b)/m
8. Solve the equation below for c, in terms of a and b.
a+4b-3c=10
3c = a+4b-10
c = (a+4b-10)/3